# Article Category

### v1

@@ 5 METHOD
We show the policy, safeguard and environment relations in Figure 1. The safeguards are applied after the policy output. To train the policy, we calculate the gradient of the reward with respect to the policy output
@@
Since the policy output depends on the previous state, full backpropagation requires unrolling the trajectory to determine how all previous policy outputs affect the current reward.
Figure 1: Description of the forward (top) and backward (bottom) pass of the provably safe analytic policy optimisation.
The following subsections describe how we safeguard the analytic policy optimisation. First, we formulate the required and desired properties of a safeguard in the aforementioned differentiable setting. Then, we introduce the two safeguards used in this work: boundary projection (BP) [BoundaryProjection] and ray mask (RM) [RayMasking].
@@ 5.1 REQUIRED AND DESIRED PROPERTIES
The required properties of a safeguard for analytic gradient-based reinforcement learning are the containment of its codomain in the safe set and the existence of a Clark generalised derivative [Clarke] everywhere, which implies that the safeguarding is at least class $C^{0}$. Beyond the required, there are additional desired properties. First, the safeguarding should be class $C^{1}$ and provide full-rank Jacobians $\frac{\partial a_{s}}{\partial a}$ everywhere to facilitate learning and prevent information loss. Second, the number of interventions by the safeguarding should be minimal throughout training and inference. Minimal interventions also serve the last desired property of fast computation. In summary, the safeguard must
@@
    be subdifferentiable everywhere, and therefore class $C^{0}$
@@
    be class $C^{1}$ and provide full-rank Jacobians everywhere, i.e. be a local diffeomorphism,
@@
The two safeguards we present offer different trade-offs between the desired properties.
@@ 5.2 BOUNDARY PROJECTION
The boundary projection safeguard [BoundaryProjection] maps actions to the safe actions with the smallest distance. It maps only the required unsafe actions to the nearest boundary point in the safe action set. We show an exemplary mapping with boundary projection from an unsafe action $a$ to a safe action $a_{s}$ in Figure 2.
Figure 2: Boundary projection maps unsafe actions to the boundary of the safe action set.
The safe action is the solution to
@@
If the safe action set is explicitly available, Equation 5 provides the containment constraint, which yields a quadratic problem.
Induced safe action sets could yield non-convex constraints, making the problem significantly more expensive. However, if the transition function is linear with respect to the action and the noise,
$\frac{\partial^{2}f}{\partial a^{2}}$ $=0$ (12)
$\frac{\partial^{2}f}{\partial w^{2}}$ $=0\,,$
the first-order Taylor expansion of the continuously differentiable transition function
$f(s_{i},a_{i},w_{i})\approx f\bigg{|}_{lin}$ $+\frac{\partial f}{\partial s}\bigg{|}_{lin}(s_{i}-s_{lin})$ (13)
$+\frac{\partial f}{\partial a}\bigg{|}_{lin}(a_{i}-a_{lin})$
$+\frac{\partial f}{\partial w}\bigg{|}_{lin}(w_{i}-w_{lin})$
around the current state is exact, if evaluated only at that state. This Taylor expansion can be utilised to compute the next set of states
$\mathcal{S}_{i+1}(s_{i},a_{i})=\left<f(s_{i},a_{i},c_{\mathcal{W}}),\frac{\partial f}{\partial w}\bigg{|}_{lin}G_{\mathcal{W}}\right>\,,$ (14)
which allows a replacement of Equation 11b with $\mathcal{S}_{i+1}(a_{i})\subseteq\mathcal{S}_{s}$. This containment constraint is either again Equation 5 for $\mathcal{W}=\emptyset$ or Equation 6 yielding again a quadratic problem.
In Euclidean space, unsafe actions are mapped orthogonally to the boundary $\partial\mathcal{A}_{s}$ of the safe action set. The distance between the initial and mapped action decreases smoothly with the distance from the unsafe action to the boundary until it is zero for safe actions. However, the mapping location can change abruptly between unsafe actions on different sides of the edges of the safe set. Moreover, any boundary point $b_{\mathcal{A}_{S}}\in\partial\mathcal{A}_{s}$ is the solution to Equation 11 for all actions
$a=b_{\mathcal{A}_{S}}+t\cdot v\quad\forall t,\,\forall v$ (15)
with the normal, unit vectors $v\in\{v\mid v\cdot(a_{s}-b_{\mathcal{A}_{S}})\leq 0,\forall a_{s}\in\mathcal{A}_{s}\}\setminus\{0\}$ and a positive scalar $t$. Intuitively, a boundary point is the nearest safe action for all actions along a normal vector. Therefore, the safeguard cannot propagate gradient components in the mapping direction, such that
@@
which is especially problematic for gradients parallel to $v$, which may happen if the optimal action is safe. In such a case, boundary projection eliminates the gradient, keeping the optimisation stuck indefinitely. It also means the Jacobian has a rank of at most $d-1$.

### v2

@@ 5 METHOD
Figure˜1 shows the general framework for provably safe, analytic gradient-based reinforcement learning. For any policy output, we apply safeguards that map the unsafe action $a_{i}$ to the safe action ${a_{s}}_{i}$. The safe action is executed in the environment, yielding the next state $s_{i+1}$ and reward $r_{i}$. To train the policy, we calculate the gradient of the reward with respect to the policy output
@@
Since the policy output and reward depend on the previous state, full backpropagation requires unrolling the trajectory to determine how all previous policy outputs affect the current reward.
Figure 1: The forward (top) pass of the provably safe policy optimisation from analytic gradients describes the integration of the safeguard in-between the policy and environment. The backward pass (bottom) visualises, how we utilise backpropagation to differentiate through the environment and safeguard to obtain the reward gradient with respect to the policy action. It also highlights how this process in principle requires the unrolling of the previous trajectory.
The following subsections detail the safeguard. First, we formulate generally required and desirable properties in the aforementioned differentiable setting. Then, we introduce the two safeguards used in this work: boundary projection (BP) [20] and ray mask (RM) [12].We structure their introduction by first explaining the principle of the safeguard, then analysing its properties, and finally present our modifications.
@@ 5.1 REQUIRED AND DESIRED PROPERTIES
A safeguard for our setting must be provably safe as described in Section˜3.2. To allow for backpropagation, it must also guarantee the existence of a Clark generalised derivative [69] everywhere, which implies that the safeguarding is at least of class $C^{0}$.
Beyond the required, there are additional desired properties. First, the safeguarding should be of class $C^{1}$ and provide full-rank Jacobians $\frac{\partial a_{s}}{\partial a}$ everywhere. A rank-deficient Jacobian can diminish the learning signal by reducing its effective dimensionality and incurring information loss. Second, the number of interventions by the safeguarding should be minimal throughout training and inference. Minimal interventions also serve the last desired property of fast computation. In summary, the safeguard must
@@
    be subdifferentiable everywhere, and therefore of class $C^{0}$
@@
    be of class $C^{1}$ and provide full-rank Jacobians everywhere, i.e. be a local diffeomorphism,
@@
In the following, we present two safeguards that offer different trade-offs between the desired properties. We summarise the properties of the safeguards in Table˜1.
Table 1: Properties of the unaltered safeguards.
Boundary Projection Ray Mask
Property˜P1 Safety
Property˜P2 Subdifferentiable
Property˜P3
Class $C^{1}$ almost everywhere almost everywhere
Jacobian rank $d-1$($d$)
Property˜P4 Interventions $\forall a\in\mathcal{A}\setminus\mathcal{A}_{s}$ $\forall a\in\mathcal{A}$
Property˜P5 Computational complexity
For $\mathcal{A}_{s}$ 1 Quadratic Problem 1 Linear Problem
For $\mathcal{S}_{s}$wedge $P_{f,lin}$ 1 Quadratic Problem (1 Conic $\vee$ 1 Quadratic) $\wedge$ 1 Linear Problem
@@ 5.2 BOUNDARY PROJECTION
The boundary projection safeguard, which was proposed in [20], maps any action to the closest safe action. By definition, it therefore only affects unsafe actions, which are mapped to the nearest boundary point in the safe action set. In Euclidean space, this corresponds to an orthogonal projection to the boundary of the safe action set. We show an exemplary mapping with boundary projection from an unsafe action $a$ to a safe action $a_{s}$ in Figure˜2.
Figure 2: Boundary projection maps unsafe actions to the boundary of the safe action set by determining the closest safe action.
The safeguard $g_{\mathcal{A}_{s},\text{BP}}(a)$ provides the safe action by solving
@@
5.2.1 PROPERTIES
The optimisation problem is always solvable given Equation˜1, and the constraint Equation˜12b satisfies Property˜P1. The implicit function theorem [49] provides the Jacobian of the solution mapping to fulfil Property˜P2.
The distance between the initial and mapped action decreases smoothly with the distance from the unsafe action to the boundary until it is zero for safe actions. However, the mapping location can change abruptly between unsafe actions on different sides of the edges of the safe set. This leads to a jump in the gradient landscape, such that it is only $C^{1}$ almost everywhere. In addition, all unsafe actions along a normal vector to the safe action set are projected to the same safe action on the boundary. Formally, any safe action on the boundary $a_{s,\partial\mathcal{A}_{s}}\in\partial\mathcal{A}_{s}$ is the solution to Equation˜12 for all actions
$a=a_{s,\partial\mathcal{A}_{s}}+t\cdot v\quad\forall t,\,\forall v$ (13)
with the normal, unit vectors $v\in\{v\mid v\cdot(a_{s}-a_{s,\partial\mathcal{A}_{s}})\leq 0,\forall a_{s}\in\mathcal{A}_{s}\}\setminus\{0\}$ and a positive scalar $t$. Therefore, the safeguard cannot propagate gradient components in the mapping direction, such that
@@ 5.2.1 PROPERTIES
which is especially problematic for gradients parallel to $v$, that occur when the optimal action is safe. In such a case, boundary projection eliminates the gradient, keeping the optimisation stuck indefinitely.